# Differentiable Pareto-Smoothed Weighting for High-Dimensional Heterogeneous Treatment Effect Estimation

### v1

@@ Abstract
There is a growing interest in estimating heterogeneous treatment effects across individuals using their high-dimensional feature attributes. Achieving high performance in such high-dimensional heterogeneous treatment effect estimation is challenging because in this setup, it is usual that some features induce sample selection bias while others do not but are predictive of potential outcomes. To avoid losing such predictive feature information, existing methods learn separate feature representations using the inverse of probability weighting (IPW). However, due to the numerically unstable IPW weights, they suffer from estimation bias under a finite sample setup. To develop a numerically robust estimator via weighted representation learning, we propose a differentiable Pareto-smoothed weighting framework that replaces extreme weight values in an end-to-end fashion. Experimental results show that by effectively correcting the weight values, our method outperforms the existing ones, including traditional weighting schemes.
@@ 1 Introduction
In this paper, we tackle the problem of estimating heterogeneous treatment effects across individuals from high-dimensional observational data. This problem, which we call high-dimensional heterogeneous treatment effect estimation, offers crucial applications. Application examples include the evaluation of medical treatment effects from numerous attributes [Curth et al., 2024, Shalit, 2020] and the assessment of the advertising effects from each user’s many attributes [Bottou et al., 2013, Wang et al., 2015].
@@
One of the fundamental difficulties in high-dimensional heterogeneous treatment effect estimation is the sample selection bias induced by confounders, i.e., the features of an individual that affect their treatment choice and outcome. For instance, in the case of medical treatment, a possible confounder is age: Older patients avoid choosing surgery due to its risk while they have higher mortality in general [Zeng et al., 2022]. Since there are fewer records of older patients who received surgery treatment, it is not easy to accurately predict the outcome of surgery on older patients, thus making it difficult to estimate the treatment effect.
@@
To address such sample selection bias, it is crucial how to break the dependence of treatment choice on confounders. A promising approach for high-dimensional setup is representation learning [Johansson et al., 2016, Shalit et al., 2017], which estimates the potential outcomes under different treatment assignments by extracting the balanced feature representation that is learned such that its distribution is identical between treated and untreated individuals.
@@
However, since this approach converts all input features to a single balanced representation, if some features are adjustment variables (a.k.a., risk factors) [Brookhart et al., 2006], which are not related to sample selection bias but are useful for outcome prediction, it may inadvertently eliminate such useful feature information, leading to inaccurate treatment effect estimation [Sauer et al., 2013]. This issue is serious, especially in high-dimensional setup where input features often contain adjustment variables. Moreover, due to the lack of prior knowledge about features, it is difficult for practitioners to correctly separate adjustment variables from confounders. Such feature separation might be impossible if one attempts to input the feature embeddings of a complex object (e.g., texts, images, and graphs) that are constructed from pre-trained generative models, including large language models (LLMs) [Keith et al., 2020].
@@
To fulfill such important but often overlooked needs, several data-driven feature separation methods have been proposed [Kuang et al., 2017, Hassanpour and Greiner, 2020, Kuang et al., 2020, Wang et al., 2023]. Among these methods, disentangled representations for counterfactual regression (DRCFR) [Hassanpour and Greiner, 2020] aim to avoid losing useful feature information for heterogeneous treatment effect estimation by learning different representations for confounders and adjustment variables. To achieve this, it employs the technique of inverse probability weighting (IPW) [Rosenbaum and Rubin, 1983], which performs weighting based on the inverse of the probability called a propensity score. However, such weights often take extreme values (especially in high-dimensional setup [Li and Fu, 2017]), and even a slight propensity score estimation error may lead to a large treatment effect estimation error.
@@
To resolve this issue, we develop a weight correction framework by utilizing the technique in extreme value statistics, called Pareto smoothing [Vehtari et al., 2024], which replaces the extreme weight values with the quantiles of generalized Pareto distribution (GPD). Indeed, Zhu et al. [2020] have already adopted Pareto smoothing and empirically shown that it can construct a numerically stable estimator of the average treatment effect (ATE) over all individuals.
@@
To estimate heterogeneous treatment effects across individuals, we propose a Pareto-smoothed weighting framework that can be combined with the weighted representation learning approach. This goal, however, is difficult to achieve because the weight correction with Pareto smoothing requires the computation of the rank (or position) of each weight value; this computation is non-differentiable and prevents gradient backpropagation. To overcome this difficulty, we utilize the technique of differentiable ranking [Blondel et al., 2020] and establish a differentiable weight correction framework founded on Pareto smoothing. This idea of combining Pareto smoothing and differentiable ranking, which have been studied in totally different fields (i.e., extreme value statistics and differentiable programming), allows for an effective end-to-end learning for high-dimensional heterogeneous treatment effect estimation.
@@
  • We propose a differentiable Pareto-smoothed weighting framework that replaces the extreme IPW weight values in an end-to-end fashion. To make this weight replacement procedure differentiable, we utilize the technique of differentiable ranking [Blondel et al., 2020].
  • Taking advantage of the differentiability, we build our weight correction framework on the neural-network-based weighted representation learning method (i.e., DR-CFR [Hassanpour and Greiner, 2020]) to perform data-driven feature separation for high-dimensional heterogeneous treatment effect estimation.
@@ 2.1 Problem Setup
Suppose that we have a sample of $n$ individuals, $\mathcal{D}=\{(a_{i},\textbf{{x}}_{i},y_{i})\}_{i=1}^{n}\overset{i.i.d.}{\sim}\operatorname{\mathrm{P}}(A,\textbf{{X}},Y)$, where $A\in\{0,1\}$ is a binary treatment ($A=1$ if an individual is treated; otherwise $A=0$), $\textbf{{X}}=[X_{1},\dots,X_{d}]^{\top}$ is $d$-dimensional features (a.k.a., covariates), and $Y$ is an outcome. Let $Y^{0}$ and $Y^{1}$ be potential outcomes i.e., the outcomes when untreated ($A=0$) and when treated ($A=1$), which are given by $Y=AY^{1}+(1-A)Y^{0}$. Then a treatment effect for an individual is defined as their difference, i.e., $Y^{1}-Y^{0}$ [Rubin, 1974].
@@
To estimate the CATE in (1), we make two standard assumptions. One is conditional ignorability, $\{Y^{0},Y^{1}\}\mathop{\perp\!\!\!\perp}A\mid\textbf{{X}}$; this conditional independence is satisfied if features X contain all confounders and include only pretreatment variables, which are not affected by treatment $A$. The other is positivity, $10<\pi(\textbf{{x}})<$ 1 for all x, where $\pi(\textbf{{x}})\coloneqq\operatorname{\mathrm{P}}(A=1\mid\textbf{{X}}=\textbf{{x}})$ is the conditional distribution model called a propensity score.
Our goal is to achieve high CATE estimation performance in high-dimensional setup, where the number of features $d$ is relatively large. Under this setup, removing the sample selection bias by transforming all features X into a single balanced representation might be overly severe, which leads to inaccurate treatment effect estimation. Although several representation learning methods aim to avoid such an overly severe balancing [Kuang et al., 2017, Wang et al., 2023], most of them are designed ATE, not CATE. By contrast, the DRCFR method [Hassanpour and Greiner, 2020] focuses on CATE estimation and effectively balances high-dimensional features X via weighted representation learning.
@@ 2.2 Weighted Representation Learning
DRCFR [Hassanpour and Greiner, 2020] is founded on the graphical model in Figure 1, where treatment $A$ is determined by functions $\Gamma(\textbf{{X}})$ and $\Delta(\textbf{{X}})$, while outcome $Y$ is given by $\Delta(\textbf{{X}})$ and $\Upsilon(\textbf{{X}})$. Following this model, it formulates $\Gamma(\textbf{{X}})$, $\Delta(\textbf{{X}})$, and $\Upsilon(\textbf{{X}})$ as three neural network encoders, each of which extracts the representation of instrumental variables (i.e., the features that influence $A$ but do not affect $Y$), confounders, and adjustment variables, respectively. For example, in medical treatment setup, a possible instrumental variable, confounder, and adjustment variable would be income, age, and smoking habits, respectively.
To learn the representations of such features, DRCFR minimizes the weighted loss by computing the weights based on the inverse of the propensity score, given by $\pi(\Gamma(\textbf{{X}}),\Delta(\textbf{{X}}))$.
@@ 3.1 Overview
To estimate the CATE in (1), following the DRCFR method [Hassanpour and Greiner, 2020], we perform weighted representation learning. To achieve this, we learn the three model components: the feature representations (i.e., $\Gamma(\textbf{{X}})$, $\Delta(\textbf{{X}})$, and $\Upsilon(\textbf{{X}})$ in Figure 1), propensity score model $\pi(\Gamma(\textbf{{X}}),\Delta(\textbf{{X}}))$, and outcome prediction models $h^{0}(\Delta(\textbf{{X}}),\Upsilon(\textbf{{X}}))$ and $h^{1}(\Delta(\textbf{{X}}),\Upsilon(\textbf{{X}}))$, where $h^{0}$ and $h^{1}$ are used to predict potential outcomes $Y^{0}$ and $Y^{1}$, respectively.
@@
First, we learn propensity score $\pi$ (while fixing other model parameters) by minimizing the cross entropy loss:
@@
where $\Omega(\cdot)$ is some regularizer that penalizes the model complexity, and $\lambda_{\pi}>0$ is a regularization parameter.
@@
where $w_{i}$ is the weight that is given by propensity score $\pi(\Gamma(\textbf{{x}}_{i}),\Delta(\textbf{{x}}_{i}))$, $l$ is the prediction loss for outcome $y_{i}$, $\lambda_{\Upsilon}>0$ and $\lambda_{-\pi}>0$ are regularization parameters, [^1] Subscript $-\pi$ denotes the other model components than $\pi$. and $\mathrm{MMD}$ denotes the kernel maximum mean discrepancy (MMD) [Gretton et al., 2012], which measures the discrepancy between empirical conditional distributions $\hat{\operatorname{\mathrm{P}}}(\Upsilon(\textbf{{X}})\mid A=0)$ and $\hat{\operatorname{\mathrm{P}}}(\Upsilon(\textbf{{X}})\mid A=1)$. Regularizing this MMD term enforces $\Upsilon$ to have no information about treatment $A$, thus making $\Upsilon(\textbf{{X}})$ a good representation of adjustment variables.
To achieve high CATE estimation performance, it is essential how to compute the weight value, i.e., $w_{i}$ in (3). In the DRCFR method [Hassanpour and Greiner, 2020], the weight is formulated using importance sampling, which employs the density-ratio-based weight to construct a weighted estimator of expected value. To estimate the expected outcome prediction loss over observed individuals ($A=a_{i}$) and the one over unobserved individuals ($A=1-a_{i}$), DRCFR formulates the weight as the sum of two density ratios:
@@
where $\pi_{a_{i}}(\textbf{{X}})=a_{i}\pi(\textbf{{X}})+(1-a_{i})(1-\pi(\textbf{{X}}))$. Here weight $w_{i}$ is proportional to the inverse of the propensity score $\pi_{a_{i}}(\Gamma(\textbf{{x}}_{i}),\Delta(\textbf{{x}}_{i}))$. Since such an IPW weight often takes an extreme value, the weight estimation is numerically unstable, leading to inaccurate CATE estimation. This issue is serious in high-dimensional setup due to the difficulty of correctly estimating the propensity score [Assaad et al., 2021].
To resolve this issue, we improve the weight stability by replacing an extreme value of $w_{i}$ in (4). To do so, we utilize the weight stabilization technique, called Pareto smoothing.
@@ 3.2 Weight Correction via Pareto Smoothing
Pareto smoothing [Vehtari et al., 2024] is the technique for improving the weight stability of importance sampling.
@@
where $L>0$ and $U>0$ are the truncation thresholds. Second, it can be combined with self-normalization, which prevents the weights from being too small or too large relative to each other by dividing each weight value by its empirical mean under the same treatment assignment:
@@
Here $\operatorname{\mathbf{I}}(a_{j}=a_{i})$ is the indicator function that takes $1$ if $a_{j}=a_{i}$; otherwise, 0. We experimentally confirmed that performing self-normalization over Pareto-smoothed weights leads to better CATE estimation performance (Section 4.1).
To construct a weighted estimator that is numerically robust to weight estimation error, Pareto smoothing replaces the extremely large weight values with the GPD quantiles. To achieve this, it takes two steps: the GPD parameter estimation and the weight replacement.
@@ 3.2.1 GPD Parameter Fitting
Suppose that a random variable $W$ follows the GPD. Then its GPD cumulative distribution function is defined as
$\operatorname{\mathrm{F}}(w)=\left\{\begin{array}[]{ll}1-\left(1+\frac{\xi(w-\mu)}{\sigma}\right)^{-\frac{1}{\xi}}&(\xi\neq 0)\\ 1-\mathrm{e}^{-\frac{w-\mu}{\sigma}}&(\xi=0)\end{array}\right.$ (12)
where $\mu\in\mathbb{R}$, $\sigma>0$, and $\xi\in\mathbb{R}$ are location, scale, and shape parameters, respectively.
We fit these GPD parameters to the $M+1$ largest IPW weight values. Here $M$ ($0<M<n$) is given by heuristic; following Vehtari et al. [2024], we determine it by
@@
where $\lfloor n\rfloor$ denotes a floor function, which returns the greatest integer that is less than or equal to $n$. Letting $w_{[1]}\leq\dots\leq w_{[n]}$ be the weights sorted in an ascending order, the $M+1$ largest ones are denoted by $w_{[n-M]},\dots,w_{[n]}$.
@@
By contrast, we estimate $\sigma$ and $\xi$ using $w_{[n-M+1]},\dots,w_{[n]}$. Among several estimators, we employ the standard one called probability weighted moment (PWM) [Hosking and Wallis, 1987],  [^2] Although we empirically observed that using PWM leads to good CATE estimation performance, GPD parameter fitting might not be easy in general. However, according to Vehtari et al. [2024, Section 6], one can effectively evaluate the reliability of GPD fitting by employing the estimated value of GPD’s shape parameter, $\hat{\xi}$, which determines the heaviness of distribution tail. which constructs the estimators of $\sigma$ and $\xi$ using the following weighted moment statistic:

### v2

@@ Abstract
There is a growing interest in estimating heterogeneous treatment effects across individuals using their high-dimensional feature attributes. Achieving high performance in such high-dimensional heterogeneous treatment effect estimation is challenging because in this setup, it is usual that some features induce sample selection bias while others do not but are predictive of potential outcomes. To avoid losing such predictive feature information, existing methods learn separate feature representations using inverse probability weighting (IPW). However, due to their numerically unstable IPW weights, these methods suffer from estimation bias under a finite sample setup. To develop a numerically robust estimator by weighted representation learning, we propose a differentiable Pareto-smoothed weighting framework that replaces extreme weight values in an end-to-end fashion. Our experimental results show that by effectively correcting the weight values, our proposed method outperforms the existing ones, including traditional weighting schemes.
@@ 1 Introduction
In this paper, we tackle the problem of estimating heterogeneous treatment effects across individuals from high-dimensional observational data. This problem, which we call high-dimensional heterogeneous treatment effect estimation, offers the following crucial applications: the evaluation of medical treatment effects from numerous attributes [Shalit, 2020] and the assessment of the advertising effects from each user’s many attributes [Bottou et al., 2013].
@@
One fundamental difficulty in high-dimensional heterogeneous treatment effect estimation is the sample selection bias induced by confounders, i.e., the features of an individual that affect their treatment choices and outcomes. For instance, in the case of medical treatment, age is a possible confounder: Elderly patients avoid choosing surgery due to its risk and generally suffer from higher mortality [Zeng et al., 2022]. Due to their treatment choice imbalance, there are often fewer records of elderly patients who have received surgical treatments. As a result, accurately predicting surgical outcomes is difficult in this age cohort, thus complicating heterogeneous treatment effect estimation.
@@
To address such sample selection bias, it is crucial to determine how to break the dependence of treatment choice on confounders. A promising approach for a high-dimensional setup is representation learning [Shalit et al., 2017], which estimates the potential outcomes under different treatment assignments by extracting the balanced feature representation that is learned such that its distribution is identical between treated and untreated individuals.
@@
However, since this approach converts all input features to a single balanced representation, if some features are adjustment variables (a.k.a., risk factors) [Brookhart et al., 2006], which are unrelated to sample selection bias but useful for outcome prediction, such useful feature information may be inadvertently eliminated, leading to inaccurate treatment effect estimation [Sauer et al., 2013]. This issue is serious, especially in high-dimensional setups where input features often contain adjustment variables. Moreover, due to the lack of prior knowledge about the features, it is difficult for practitioners to correctly separate adjustment variables from confounders. Such feature separation might be impossible if one attempts to input the feature embeddings of a complex object (e.g., texts, images, and graphs) that are constructed from pre-trained generative models, including large language models (LLMs) [Keith et al., 2020].
@@
To fulfill such important but often overlooked needs, several data-driven feature separation methods have been proposed [Kuang et al., 2017, Hassanpour and Greiner, 2020, Kuang et al., 2020, Wang et al., 2023]. Among them, disentangled representations for counterfactual regression (DRCFR) [Hassanpour and Greiner, 2020] aims to avoid losing useful feature information for heterogeneous treatment effect estimation by learning different representations for confounders and adjustment variables. To achieve this, this method employs a technique of inverse probability weighting (IPW) [Rosenbaum and Rubin, 1983], which performs weighting based on the inverse of the probability called a propensity score. However, such weights often take extreme values (especially in high-dimensional setups [Li and Fu, 2017]), and even alight propensity score estimation error may lead to large treatment effect estimation error.
@@
To resolve this issue, we develop a weight correction framework that utilizes a technique in extreme value statistics, called Pareto smoothing [vehtari2015pareto], which replaces the extreme weight values with the quantiles of generalized Pareto distribution (GPD). Indeed, Zhu et al. [2020] already adopted Pareto smoothing and empirically showed that it can construct a numerically stable estimator of the average treatment effect (ATE) over all individuals.
@@
To estimate heterogeneous treatment effects across individuals, we propose a Pareto-smoothed weighting framework that can be combined with the weighted representation learning approach. However, achieving this goal is difficult because weight correction with Pareto smoothing requires the computation of the rank (position) of each weight value; this computation is non-differentiable and prevents gradient backpropagation. To overcome this difficulty, we utilize a differentiable ranking technique [Blondel et al., 2020] and establish a differentiable weight correction framework founded on Pareto smoothing. This idea of combining Pareto smoothing and differentiable ranking, both of which have been studied in completely different fields (i.e., extreme value statistics and differentiable programming), allows for effective end-to-end learning for high-dimensional heterogeneous treatment effect estimation.
@@
  • We propose a differentiable Pareto-smoothed weighting framework that replaces extreme IPW weight values in an end-to-end fashion. To make this weight replacement procedure differentiable, we utilize a differentiable ranking technique [Blondel et al., 2020].
  • Exploiting differentiability, we build our weight correction framework on the neural-network-based weighted representation learning method (i.e., DR-CFR [Hassanpour and Greiner, 2020]) and perform data-driven feature separation for high-dimensional heterogeneous treatment effect estimation.
@@ 2.1 Problem Setup
Suppose we have a sample of $n$ individuals, $\mathcal{D}=\{(a_{i},\textbf{{x}}_{i},y_{i})\}_{i=1}^{n}\overset{i.i.d.}{\sim}\operatorname{\mathrm{P}}(A,\textbf{{X}},Y)$, where $A\in\{0,1\}$ is a binary treatment ($A=1$ if an individual is treated; otherwise $A=0$), $\textbf{{X}}=[X_{1},\dots,X_{d}]^{\top}$ is the $d$-dimensional features (a.k.a., covariates), and $Y$ is an outcome. Let $Y^{0}$ and $Y^{1}$ be potential outcomes i.e., the outcomes when untreated ($A=0$) and when treated ($A=1$), which are given by $Y=AY^{1}+(1-A)Y^{0}$. A treatment effect for an individual is defined as their differences, i.e., $Y^{1}-Y^{0}$ [Rubin, 1974].
@@
To estimate the CATE in (1), we make two standard assumptions. One is conditional ignorability, $\{Y^{0},Y^{1}\}\mathop{\perp\!\!\!\perp}A\mid\textbf{{X}}$; this conditional independence is satisfied if features X contain all of the confounders and include only pretreatment variables, which are not affected by treatment $A$. The other is positivity, $10<\pi(\textbf{{x}})<$ 1 for all x, where $\pi(\textbf{{x}})\coloneqq\operatorname{\mathrm{P}}(A=1\mid\textbf{{X}}=\textbf{{x}})$ is a conditional distribution model called a propensity score.
Our goal is to achieve a high CATE estimation performance in a high-dimensional setup, where the number of features $d$ is relatively large. Under this setup, removing the sample selection bias by transforming all features X into a single balanced representation might be overly severe, which leads to inaccurate treatment effect estimation. Although several representation learning methods aim to avoid such an overly severe balancing [Kuang et al., 2017, Wang et al., 2023], most are designed for ATE, not for CATE. By contrast, the DRCFR method [Hassanpour and Greiner, 2020] focuses on CATE estimation and effectively balances high-dimensional features X by weighted representation learning.
@@ 2.2 Weighted Representation Learning
DRCFR [Hassanpour and Greiner, 2020] is based on the graphical model in Figure 1, where treatment $A$ is determined by functions $\Gamma(\textbf{{X}})$ and $\Delta(\textbf{{X}})$, while outcome $Y$ is given by $\Delta(\textbf{{X}})$ and $\Upsilon(\textbf{{X}})$. Following this model, $\Gamma(\textbf{{X}})$, $\Delta(\textbf{{X}})$, and $\Upsilon(\textbf{{X}})$ are formulated as three neural network encoders, each of which extracts the representation of instrumental variables (i.e., the features that influence $A$ but not $Y$), confounders, and adjustment variables, respectively. For example, in medical treatment setups, a possible instrumental variable, confounder, and adjustment variable might be income, age, and smoking habits, respectively.
To learn the representations of such features, DRCFR minimizes the weighted loss by computing the weights based on the inverse of propensity scores, given by $\pi(\Gamma(\textbf{{X}}),\Delta(\textbf{{X}}))$.
@@ 3.1 Overview
To estimate the CATE in (1), following DRCFR [Hassanpour and Greiner, 2020], we perform weighted representation learning. We learn three model components: the feature representations (i.e., $\Gamma(\textbf{{X}})$, $\Delta(\textbf{{X}})$, and $\Upsilon(\textbf{{X}})$ in Figure 1), propensity score model $\pi(\Gamma(\textbf{{X}}),\Delta(\textbf{{X}}))$, and outcome prediction models $h^{0}(\Delta(\textbf{{X}}),\Upsilon(\textbf{{X}}))$ and $h^{1}(\Delta(\textbf{{X}}),\Upsilon(\textbf{{X}}))$, where $h^{0}$ and $h^{1}$ are used to predict potential outcomes $Y^{0}$ and $Y^{1}$.
@@
First, we learn propensity score $\pi$ (while the fixing other model parameters) by minimizing the cross entropy loss:
@@
where $\Omega(\cdot)$ is a regularizer that penalizes the model complexity, and $\lambda_{\pi}>0$ is a regularization parameter.
@@
[…]
To achieve a high CATE estimation performance, how to compute the weight value (i.e., $w_{i}$ in (3)) is essential. In the DRCFR method [Hassanpour and Greiner, 2020], weight is formulated using importance sampling, which employs a density-ratio-based weight to construct a weighted estimator of expected value. To estimate the expected outcome prediction losses over the observed individuals ($A=a_{i}$) and the unobserved individuals ($A=1-a_{i}$), DRCFR formulates the weight as the sum of two density ratios:
@@
where $\pi_{a_{i}}(\textbf{{X}})=a_{i}\pi(\textbf{{X}})+(1-a_{i})(1-\pi(\textbf{{X}}))$. Here weight $w_{i}$ is proportional to the inverse of propensity score $\pi_{a_{i}}(\Gamma(\textbf{{x}}_{i}),\Delta(\textbf{{x}}_{i}))$. Since such an IPW weight often takes an extreme value, the weight estimation is numerically unstable, leading to inaccurate CATE estimation. This issue is serious in a high-dimensional setup due to the difficulty of correctly estimating propensity scores [Assaad et al., 2021].
To resolve this issue, we improve the weight stability by replacing an extreme value of $w_{i}$ in Eq. (4). with a weight stabilization technique, called Pareto smoothing.
@@ 3.2 Weight Correction via Pareto Smoothing
Pareto smoothing [Vehtari et al., 2024] is a technique for improving the weight stability of importance sampling.
@@
where $L>0$ and $U>0$ are the truncation thresholds. Second, it can be combined with self-normalization, which prevents the weights from being too small or too large relative to each other by dividing each weight value by its empirical mean under identical treatment assignment:
@@
Here $\operatorname{\mathbf{I}}(a_{j}=a_{i})$ is an indicator function that takes $1$ if $a_{j}=a_{i}$; otherwise, 0. We experimentally confirmed that performing self-normalization over Pareto-smoothed weights leads to better CATE estimation performance (Section 4.1).
To construct a weighted estimator that is numerically robust to weight estimation error, Pareto smoothing replaces the extremely large weight values with GPD quantiles in two steps: GPD parameter estimation and weight replacement.
@@ 3.2.1 GPD Parameter Fitting
Suppose that random variable $W$ follows the GPD. Then its GPD cumulative distribution function is defined as
$\operatorname{\mathrm{F}}(w)=\left\{\begin{array}[]{ll}1-\left(1+\frac{\xi(w-\mu)}{\sigma}\right)^{-\frac{1}{\xi}}&(\xi\neq 0),\\ 1-\mathrm{e}^{-\frac{w-\mu}{\sigma}}&(\xi=0)\end{array}\right.$ (12)
where $\mu\in\mathbb{R}$, $\sigma>0$, and $\xi\in\mathbb{R}$ are location, scale, and shape parameters.
We fit these GPD parameters to the $M+1$ largest IPW weight values. Here $M$ ($0<M<n$) is given by heuristics; following Vehtari et al. [2024], we determine it by
@@
where $\lfloor n\rfloor$ denotes a floor function, which returns the greatest integer that is less than or equal to $n$. Letting $w_{[1]}\leq\dots\leq w_{[n]}$ be the weights sorted in ascending order, the $M+1$ largest ones are denoted by $w_{[n-M]},\dots,w_{[n]}$.
@@
By contrast, we estimate $\sigma$ and $\xi$ using $w_{[n-M+1]},\dots,w_{[n]}$. Among several estimators, we employ the standard method termed the probability weighted moment (PWM) [Hosking and Wallis, 1987],  [^2] Although we empirically observed that using PWM leads to good CATE estimation performance, GPD parameter fitting might not be easy in general. However, according to Vehtari et al. [2024, Section 6], one can evaluate the reliability of GPD fitting by employing the estimated value of GPD’s shape parameter, $\hat{\xi}$, which determines the heaviness of the distribution tail. which constructs the estimators of $\sigma$ and $\xi$ using the following weighted moment statistic: